CaiCai:

If s = 1 over 3 unit and A = 18s2, what is the value of A, in square units? ____ square unit(s)

Yggdrasil:

is the formula like this? \[A=18s ^{2}\]

CaiCai:

No it is not

Yggdrasil:

so is it 18 times s, times 2?

CaiCai:

yes... i think i am no help

Yggdrasil:

what does "over 3 unit" mean?

CaiCai:

It means 1/3

Yggdrasil:

s = 1/3? A = 18*1/3*2 = 13 square units

CaiCai:

Thank you

Yggdrasil:

Let's just pretend I understood the question